Solve system by substitution or addition, whichever is easier.
2y - x = 3
x = 3y - 5

Answer by algebrapro18(206)

Since one of the equations has a variable solved for already, lets use substitution.

The steps in substitution are as follows:

1) Solve one of the equations for a variable.
2) Plug in the value of that variable into the other equation and solve for the one variable left.
3) Plug that value into 1 and solve for the first variable.

Since step one has already been done for us we can just jump to step to.

2y - x = 3
x = 3y - 5

Plugging x = 3y-5 into the top equation you get:

2y -(3y-5) = 3

now we just solve for y

2y -(3y-5) = 3
2y - 3y + 5 = 3
-y + 5 = 3
-y = -2
y = 2

Now we move on to step 3. Plugging the value y = 2 into the second equation we get:

x = 3(2)-5
x = 6 - 5
x = 1

now we are done, our solution is (1,2) or x = 1 and y =2 depending on what form we want to leave it in.